# Parc Científic UV, c/ Catedrático José Beltrán, 2, E-46980 Paterna (València), Spain

### v2

@@ 3.1.1 Tree level: the Seesaw case
where $m_{\rm R}$ is a $n_{R}\times n_{R}$ complex Majorana mass matrix which, without loss of generality, can be chosen to be diagonal with positive eigenvalues, and $Y$ is a $n_{L}\times n_{R}$ complex Yukawa matrix. Here $L=(\nu_{L},e_{L})^{T}$ are the left-handed lepton doublets. [^5] This flavor structure also applies, for instance, to Type-III seesaw and Model $A_{1}$ in Ref. Giarnetti:2023osf . We denote by $H$ the SM Higgs doublet, which in the unitary gauge after electroweak symmetry breaking is given by
@@
$m_{\text{light}}\simeq m_{D}m_{R}^{-1}m_{D}^{T}\in\mathbb{C}^{n_{L}\times n_{L}},\quad m_{\text{heavy}}\simeq m_{R}\in\mathbb{C}^{n_{R}\times n_{R}}.$ (19)
@@
where $Y_{1}=Y^{T}$ and $M_{11}=v^{2}/2\,m^{-1}_{R}$. Therefore, it has the form of Eq. (3) with $M_{12}=M_{21}=M_{22}=0$, and the usual CI parametrisation Casas:2001sr can be used to express the Yukawa $Y_{1}$ in terms of light neutrino observables and other unknown free parameters. The most general neutrino Yukawa coupling compatible with low energy data, in the basis where the charged leptons Yukawa couplings and the $M_{11}$ are diagonal, reads:
$Y_{1}$ $=D^{-1/2}_{M_{11}}\,R_{\ell}\,D_{\sqrt{m}}\,U^{\dagger}\,$ (21)
$=\frac{\sqrt{2}}{v}\,m_{R}^{1/2}\,R_{\ell}\,D_{\sqrt{m}}\,U^{\dagger}\,,\qquad\text{Standard\,\,CI\,\, parametrisation},$ (22)
where $D^{-1/2}_{M_{11}}$ is the diagonal matrix of the square roots of the eigenvalues of $M_{11}$, $D_{\sqrt{m}}$ is the diagonal matrix of roots of the physical masses $m_{i}$ of the light neutrinos, and $R_{\ell}$ is the $n_{R}\times n_{L}$ orthogonal matrix that includes all the free parameters describing the neutrino sector. In Sec. 5, we discuss a rewriting of the general neutrino mass matrix in Eq. (18), which could potentially serve to obtain a GCI parametrisation outside of the seesaw limit, or one that includes higher order corrections. We also show in App. A how the standard CI may be derived in the seesaw limit using this approach.
Several other models exist in the literature where only the $M_{11}$ element is different from zero at tree level, like Type-III seesaw Foot:1988aq , Model $A_{1}$ in Ref. Giarnetti:2023osf , and other variants of the Type-I seesaw Branco:2020yvs ; CentellesChulia:2024uzv ; Grimus:2001ex .
3.1.2 Loop level: the Scotogenic Model
The Scotogenic Model (ScM) Tao96 ; Ma:2006km is based on a $\mathbb{Z}_{2}$ symmetry, see also Ref. Ma:1998dn . The particle content with the associated charges can be found in Table 3. The relevant terms in the Lagrangian of the model are:
@@ 3.1.2 Loop level: the Scotogenic Model
$\mathcal{V}_{\rm ScM}=$ $\,-\,m_{H}^{2}H^{\dagger}H\,+\,\lambda_{H}(H^{\dagger}H)^{2}\,+\,m_{\Phi}^{2}\Phi^{\dagger}\Phi\,+\,\lambda_{\Phi}(\Phi^{\dagger}\Phi)^{2}\,$ (24)
$\,+\,\lambda_{H\Phi}(H^{\dagger}H)(\Phi^{\dagger}\Phi)\,+\,\lambda_{H\Phi,2}(H^{\dagger}\Phi)(\Phi^{\dagger}H)\,+\,\frac{\lambda_{H\Phi,3}}{2}\left[(H^{\dagger}\Phi)^{2}\,+\,{\rm H.c.}\right]\,.$
The scalar mass spectrum, derived from the potential in Eq. (24), reads
@@
$m_{\phi_{0}^{I}}$ $=$ $\sqrt{m_{\Phi}^{2}\,+\,\frac{1}{2}\left(\lambda_{H\Phi}\,+\,\lambda_{H\Phi,2}\,-\,\lambda_{H\Phi,3}\right)v^{2}}\,$, (26)
$m_{\phi^{+}}$ $=$ $\sqrt{m_{\Phi}^{2}\,+\,\frac{1}{2}\lambda_{H\Phi}\,v^{2}}\,.$ (27)
Here $\phi_{0}^{R}$ and $\phi_{0}^{I}$ denote the real and imaginary parts of the neutral field $\phi_{0}$ in the additional scalar doublet, and $\phi^{+}$ its charged component. The scalars $\phi_{0}^{R}$ and $\phi_{0}^{I}$ acquire a mass splitting proportional to the quartic coupling $\lambda_{H\Phi,3}$; taking $\lambda_{H\Phi,3}$ small is technically natural in the ’t Hooft sense tHooft:1979rat .
@@
Therefore, the seesaw-like flavor structure allows to use the standard CI parametrisation Casas:2001sr , after trivially taking into account the different global factor. This result is in agreement with that found in Ref. Ibarra:2016dlb , where the ScM was studied in the context of direct detection of Dark Matter.
At one loop, another similar example is the ScotoSinglet Model Beniwal:2020hjc ; see Refs. Cacciapaglia:2020psm ; Bonnet:2012kz ; Restrepo:2013aga for other one-loop models and Ref. Cai:2017jrq for a review of radiative neutrino mass models.
3.2 Off-diagonal mass matrix
In the following we consider the scenario where only the crossed terms are non-vanishing, i.e., $M_{11}=M_{22}=0$. We want to show how the CI parametrisation can be implemented for such a broad type of scenario, without considering a particular model; this means that we do not consider if the neutrino mass matrix leads to a massless neutrino eigenvalue or to a specific mass ordering. We will go through some other examples in the following sections. The light neutrino mass matrix reads of Eq. (3) reads
@@ 3.2 Off-diagonal mass matrix
In standard seesaw models, the mixing of the charged leptons typically scales as $\Theta\sim\sqrt{m_{\nu}/M_{N}}$, with $M_{N}$ the mass of the heavy leptons that leads to a suppression of the size of the Yukawa couplings, unless special mass textures are imposed in accordance with lepton number conservation WYLER1983205 ; Mohapatra:1986aw ; Mohapatra:1986bd ; BERNABEU1987303 ; Branco:1988ex ; Barr:2003nn ; Kersten:2007vk ; Gavela:2009cd ; Hernandez:2022ivz . This feature is shared by both the Linear seesaw and the Generalised-Scotogenic Models, discussed below as examples of tree-level and loop-level neutrino mass generation, respectively. In the Linear seesaw scenario, lepton number violation (LNV) arises only through a small perturbation in the mass matrix, while in the GScM non-zero tree level neutrino masses are protected by an additional $U(1)_{\text{DM}}$ symmetry. As a consequence, in both models the violation of lepton number is strongly suppressed, which naturally leads to a hierarchical Yukawa structure, i.e., $Y_{2}\ll Y_{1}$. Therefore, within the CI parametrisation, it is immediate to take the limit of large imaginary parts of the complex angle zz, [^7] In the large imaginary part limit of the angle $z$, the imaginary component is assumed to be positive. corresponding to this strong hierarchy, which yields a rich phenomenology. For Inverted Ordering (IO), we obtain:
${Y_{1}}=\dfrac{e^{-\text{Im}z}}{\sqrt{2\overline{M}_{1}}}\Big[u_{1}^{\ast}\sqrt{m_{1}}\mp i\,u_{2}^{\ast}\sqrt{m_{2}}\Big]\,,\quad Y_{2}=\dfrac{e^{\text{Im}z}}{\sqrt{2\overline{M}_{1}}}\Big[u_{1}^{\ast}\sqrt{m_{1}}\pm i\,u_{2}^{\ast}\sqrt{m_{2}}\Big]\,$, (35)
@@
${Y_{1}}=\dfrac{e^{-\text{Im}z}}{\sqrt{2\overline{M}_{1}}}\Big[u_{2}^{\ast}\sqrt{m_{2}}\mp i\,u_{3}^{\ast}\sqrt{m_{3}}\Big]\,,\quad Y_{2}=\dfrac{e^{\text{Im}z}}{\sqrt{2\overline{M}_{1}}}\Big[u_{2}^{\ast}\sqrt{m_{2}}\pm i\,u_{3}^{\ast}\sqrt{m_{3}}\Big]\,$, (36)
where $u_{i}$ are the two columns of the PMNS mixing matrix. It is worth to notice that this result is the large mixing angle limit of the parametrisation of Ref. Ibarra:2003up , obtained for the seesaw scenario with two right-handed neutrinos. Also, a similar parametrisation can be found in Ref. Ibarra:2011xn in the context of low scale seesaw with two right-handed neutrinos.
@@
3.2.1 Tree level: the Linear Seesaw Model
The Linear seesaw (LSS) was first introduced in Refs. Akhmedov:1995ip ; Akhmedov:1995vm for left-right symmetric models and also in grand unified theories Barr:2003nn ; Malinsky:2005bi . Variations of the neutrino mass matrix have been proposed in Refs. Abada:2007ux ; Gavela:2009cd . More recently, models with new scalars in higher $SU(2)$ representations have been proposed Giarnetti:2023osf ; Giarnetti:2023dcr . In general, the model consists of $n_{1}$ right-handed neutrinos $N$ and $n_{2}$ singlet fermions $S$, so that the relevant terms in the Lagrangian are:
@@ 3.2.1 Tree level: the Linear Seesaw Model
LLLLHHHHNNSS
@@
One may directly use the GCI parametrisation, Eq. (11), with $V$ given by Eq. (33). For the $n_{1}=n_{2}=1$ case, the simpler forms for the Yukawas provided in Eqs. (35) and (36) may be used for IO and NO, respectively. Note that $Y_{S}$ violates lepton number in two units and therefore it is small in ’t Hooft sense tHooft:1979rat . This correspond to the large imaginary part of the complex angle $z$. This flavor structure also applies to Models $A_{2}$ and $B_{i}$ in Ref. Giarnetti:2023osf . The parametrisation presented for the Linear seesaw model can be mapped onto the parametrisation used in Ref. Hernandez:2022ivz in the case of degenerate heavy neutrinos by taking $\text{Re[z]}\rightarrow\theta/2$. Note that in that reference the Majorana phase is shifted by $\pi/2$ in order to provide all positive neutrino masses, whereas here the result is automatically obtained.
3.2.2 Loop level: the Generalised Scotogenic Model
Here we discuss the Generalised Scotogenic Model (GScM) Hagedorn:2018spx , a generalisation of the ScM Model Tao96 ; Ma:2006km based on a global $\text{U}(1)_{\text{DM}}$ symmetry instead of the $\mathbb{Z}_{2}$ symmetry of ScM. We follow closely the notation of Ref. Hagedorn:2018spx . The SM is augmented by two additional scalar doublets and at least one vector-like Dirac fermion, all charged under the $\text{U}(1)_{\text{DM}}$ symmetry. The particle content with the associated charges can be found in Table 2. The relevant terms in the Lagrangian are
@@ 3.2.2 Loop level: the Generalised Scotogenic Model
$\mathcal{V}_{\rm GScM}$ $=-m_{H}^{2}H^{\dagger}H+\lambda_{H}(H^{\dagger}H)^{2}+m_{\Phi}^{2}\Phi^{\dagger}\Phi+\lambda_{\Phi}(\Phi^{\dagger}\Phi)^{2}+m_{\Phi^{\prime}}^{2}\Phi^{\prime\dagger}\Phi^{\prime}+\lambda_{\Phi^{\prime}}(\Phi^{\prime\dagger}\Phi^{\prime})^{2}$
$\quad+\lambda_{H\Phi}(H^{\dagger}H)(\Phi^{\dagger}\Phi)+\lambda_{H\Phi^{\prime}}(H^{\dagger}H)(\Phi^{\prime\dagger}\Phi^{\prime})+\lambda_{\Phi\Phi^{\prime}}(\Phi^{\dagger}\Phi)(\Phi^{\prime\dagger}\Phi^{\prime})$
$\quad+\lambda_{H\Phi,2}(H^{\dagger}\Phi)(\Phi^{\dagger}H)+\lambda_{H\Phi^{\prime},2}(H^{\dagger}\tilde{\Phi}^{\prime})(\tilde{\Phi}^{\prime\dagger}H)+\lambda_{\Phi\Phi^{\prime},2}(\Phi^{\dagger}\tilde{\Phi}^{\prime})(\tilde{\Phi}^{\prime\dagger}\Phi)$
$\quad+\lambda_{H\Phi\Phi^{\prime}}\left[(H^{\dagger}\tilde{\Phi}^{\prime})(H^{\dagger}\Phi)+\text{H.c.}\right].$ (45)
@@
[…]
$\Phi\equiv(\phi^{+},\phi_{0})^{T}$ 1 2 $\phantom{-}1/2$ 1
$\Phi^{\prime}\equiv({\phi_{0}}^{\prime},{\phi^{-}}^{\prime})^{T}$ 1 2 1/2-1/2 1
Vector-like fermion $\psi$ 1 1 $\phantom{-}0$ 1
[…]
@@
$(m_{\nu})_{\alpha\beta}=\dfrac{\sin 2\theta}{32\pi^{2}}\sum_{i}\left(Y_{\alpha i}Y^{\prime}_{\beta i}M_{i}\,+Y^{\prime}_{\alpha i}Y_{\beta i}M_{i}\,\right)\mathcal{F}(m_{\eta_{0}},m_{\eta_{0}^{\prime}},M_{i})\,$, (51)
@@
$Y_{1}=Y\,,\qquad Y_{2}=Y^{\prime}\,,\qquad M_{12}=\dfrac{\sin 2\theta}{32\pi^{2}}\,{\rm diag}(M_{i}\,\mathcal{F}(M_{i}))\,.$ (52)
For the $n_{1}=n_{2}=1$ case, Eqs. (35) and (36) may be used for IO and NO, respectively. As a consistency check, we note that the parametrisation derived from the GCI construction coincides with that employed in Refs. Cai:2014kra ; Hagedorn:2018spx by replacing $\zeta\rightarrow e^{-\text{Im}z}$. For suitable lepton number charge assignments of the additional scalar and fermion fields, lepton number violation originates from the Yukawa coupling $Y$, which can thus be regarded as naturally small in the sense of ’t Hooft naturalness. This corresponds to the large imaginary part of the complex angle $z$.
@@
3.3 Diagonal plus off-diagonal mass matrix

### v3

@@ III.1.1 Tree level: the Seesaw case
where $m_{\rm R}$ is a $n_{R}\times n_{R}$ complex Majorana mass matrix which, without loss of generality, can be chosen to be diagonal with positive eigenvalues, and $Y$ is a $n_{L}\times n_{R}$ complex Yukawa matrix. Here $L=(\nu_{L},e_{L})^{T}$ are the left-handed lepton doublets. [^5] This flavour structure also applies, for instance, to Type-III seesaw and Model $A_{1}$ in Ref. [47]. We denote by $H$ the SM Higgs doublet, which in the unitary gauge after electroweak symmetry breaking is given by
@@
$m_{\text{light}}\simeq m_{D}m_{R}^{-1}m_{D}^{T}\in\mathbb{C}^{n_{L}\times n_{L}},\,$
$m_{\text{heavy}}\simeq m_{R}\in\mathbb{C}^{n_{R}\times n_{R}}.$ (20)
@@
where $Y_{1}=Y^{T}$ and $M_{11}=v^{2}/2\,m^{-1}_{R}$. Therefore, it has the form of Eq. (II) with $M_{12}=M_{21}=M_{22}=0$, and the usual CI parametrisation [24] can be used to express the Yukawa $Y_{1}$ in terms of light neutrino observables and other unknown free parameters. The most general neutrino Yukawa coupling compatible with low energy data, in the basis where the charged leptons Yukawa couplings and the $M_{11}$ are diagonal, reads:
$Y_{1}$ $=D^{-1/2}_{M_{11}}\,R_{\ell}\,D_{\sqrt{m}}\,U^{\dagger}\,$
$=\frac{\sqrt{2}}{v}\,m_{R}^{1/2}\,R_{\ell}\,D_{\sqrt{m}}\,U^{\dagger}\,$ (22)
$\text{Standard CI parametrisation},$
where $D^{-1/2}_{M_{11}}$ is the diagonal matrix of the square roots of the eigenvalues of $M_{11}$, $D_{\sqrt{m}}$ is the diagonal matrix of roots of the physical masses $m_{i}$ of the light neutrinos, and $R_{\ell}$ is the $n_{R}\times n_{L}$ complex orthogonal matrix that includes all the free parameters describing the neutrino sector. In Sec. V, we discuss a rewriting of the general neutrino mass matrix in Eq. (19), which could potentially serve to obtain a GCI parametrisation outside of the seesaw limit, or one that includes higher order corrections. We also show in App. A how the standard CI may be derived in the seesaw limit using this approach.
Several other models exist in the literature where only the $M_{11}$ element is different from zero at tree level, like Type-III seesaw [43], Model $A_{1}$ in Ref. [47], and other variants of the Type-I seesaw [19, 25, 51].
III.1.2 Loop level: the Scotogenic Model
The Scotogenic Model (ScM) [93, 72] is based on a $\mathbb{Z}_{2}$ symmetry, see also Ref. [71]. The particle content with the associated charges can be found in Table 1. The relevant terms in the Lagrangian of the model are:
@@ III.1.2 Loop level: the Scotogenic Model
$\mathcal{V}_{\rm ScM}=$ $\,-\,m_{H}^{2}H^{\dagger}H+\lambda_{H}(H^{\dagger}H)^{2}\,+\,m_{\Phi}^{2}\Phi^{\dagger}\Phi\,+\,\lambda_{\Phi}(\Phi^{\dagger}\Phi)^{2}\,$
$+\,\lambda_{H\Phi}(H^{\dagger}H)(\Phi^{\dagger}\Phi)+\lambda_{H\Phi,2}(H^{\dagger}\Phi)(\Phi^{\dagger}H)\,$
$+\,\frac{\lambda_{H\Phi,3}}{2}\left[(H^{\dagger}\Phi)^{2}\,+\,{\rm H.c.}\right]\,.$ (24)
The scalar mass spectrum, derived from the potential in Eq. (III.1.2), reads
@@
$m_{\phi_{0}^{I}}$ $=\sqrt{m_{\Phi}^{2}\,+\,\frac{1}{2}\left(\lambda_{H\Phi}\,+\,\lambda_{H\Phi,2}\,-\,\lambda_{H\Phi,3}\right)v^{2}}\,$,
$m_{\phi^{+}}$ $=\sqrt{m_{\Phi}^{2}\,+\,\frac{1}{2}\lambda_{H\Phi}\,v^{2}}\,.$
Here $\phi_{0}^{R}$ and $\phi_{0}^{I}$ denote the real and imaginary parts of the neutral field $\phi_{0}$ in the additional scalar doublet, and $\phi^{+}$ its charged component. The scalars $\phi_{0}^{R}$ and $\phi_{0}^{I}$ acquire a mass splitting proportional to the quartic coupling $\lambda_{H\Phi,3}$; taking $\lambda_{H\Phi,3}$ small is technically natural in the ’t Hooft sense [1].
@@
Therefore, the seesaw-like flavour structure allows to use the standard CI parametrisation [24], after trivially taking into account the different global factor. This result is in agreement with that found in Ref. [64], where the ScM was studied in the context of direct detection of Dark Matter.
At one loop, another similar example is the ScotoSinglet Model [14]; see Refs. [21, 17, 90] for other one-loop models and Ref. [23] for a review of radiative neutrino mass models.
III.2 Off-diagonal mass matrix
In the following we consider the scenario where only the crossed terms are non-vanishing, i.e., $M_{11}=M_{22}=0$. We want to show how the CI parametrisation can be implemented for such a broad type of scenario, without considering a particular model; this means that we do not consider if the neutrino mass matrix leads to a massless neutrino eigenvalue or to a specific mass ordering. We will go through some other examples in the following sections. The light neutrino mass matrix reads of Eq. (II) reads
@@ III.2 Off-diagonal mass matrix
In standard seesaw models, the mixing of the charged leptons typically scales as $\Theta\sim\sqrt{m_{\nu}/M_{N}}$, with $M_{N}$ the mass of the heavy leptons that leads to a suppression of the size of the Yukawa couplings, unless special mass textures are imposed in accordance with lepton number conservation [96, 80, 79, 15, 18, 12, 66, 45, 57]. This feature is shared by both the Linear seesaw and the Generalised-Scotogenic Models, discussed below as examples of tree-level and loop-level neutrino mass generation, respectively. In the Linear seesaw scenario, lepton number violation (LNV) arises only through a small perturbation in the mass matrix, while in the GScM non-zero tree level neutrino masses are protected by an additional $U(1)_{\text{DM}}$ symmetry. As a consequence, in both models the violation of lepton number is strongly suppressed, which naturally leads to a hierarchical Yukawa structure, i.e., $Y_{2}\ll Y_{1}$. Therefore, within the CI parametrisation, it is immediate to take the limit of large imaginary parts of the complex angle zz, [^7] In the large imaginary part limit of the angle $z$, the imaginary component is assumed to be positive. corresponding to this strong hierarchy, which yields a rich phenomenology. For Inverted Ordering (IO), we obtain:
${Y_{1}}=\dfrac{e^{-\text{Im}z}}{\sqrt{2\overline{M}_{1}}}\Big[u_{1}^{\ast}\sqrt{m_{1}}\mp i\,u_{2}^{\ast}\sqrt{m_{2}}\Big]\,$,
$Y_{2}=\dfrac{e^{\text{Im}z}}{\sqrt{2\overline{M}_{1}}}\Big[u_{1}^{\ast}\sqrt{m_{1}}\pm i\,u_{2}^{\ast}\sqrt{m_{2}}\Big]\,$, (33)
@@
${Y_{1}}=\dfrac{e^{-\text{Im}z}}{\sqrt{2\overline{M}_{1}}}\Big[u_{2}^{\ast}\sqrt{m_{2}}\mp i\,u_{3}^{\ast}\sqrt{m_{3}}\Big]\,$,
$Y_{2}=\dfrac{e^{\text{Im}z}}{\sqrt{2\overline{M}_{1}}}\Big[u_{2}^{\ast}\sqrt{m_{2}}\pm i\,u_{3}^{\ast}\sqrt{m_{3}}\Big]\,$, (34)
where $u_{i}$ are the two columns of the PMNS mixing matrix. It is worth to notice that this result is the large mixing angle limit of the parametrisation of Ref. [63], obtained for the seesaw scenario with two right-handed neutrinos. Also, a similar parametrisation can be found in Ref. [61] in the context of low scale seesaw with two right-handed neutrinos.
@@
LLLLHHHHNNSS
[…]
III.2.1 Tree level: the Linear Seesaw Model
The Linear seesaw (LSS) was first introduced in Refs. [5, 4] for left-right symmetric models and also in grand unified theories [12, 76]. Variations of the neutrino mass matrix have been proposed in Refs. [2, 45]. More recently, models with new scalars in higher $SU(2)$ representations have been proposed [47, 48]. In general, the model consists of $n_{1}$ right-handed neutrinos $N$ and $n_{2}$ singlet fermions $S$, so that the relevant terms in the Lagrangian are:
@@ III.2.1 Tree level: the Linear Seesaw Model
[…]
$\Phi\equiv(\phi^{+},\phi_{0})^{T}$ 1 2 $\phantom{-}1/2$ 1
$\Phi^{\prime}\equiv({\phi_{0}}^{\prime},{\phi^{-}}^{\prime})^{T}$ 1 2 1/2-1/2 1
Vector-like fermion $\psi$ 1 1 $\phantom{-}0$ 1
@@
One may directly use the GCI parametrisation, Eq. (11), with $V$ given by Eq. (31). For the $n_{1}=n_{2}=1$ case, the simpler forms for the Yukawas provided in Eqs. (III.2) and (III.2) may be used for IO and NO, respectively. Note that $Y_{S}$ violates lepton number in two units and therefore it is small in ’t Hooft sense [1]. This correspond to the large imaginary part of the complex angle $z$. This flavour structure also applies to Models $A_{2}$ and $B_{i}$ in Ref. [47]. The parametrisation presented for the Linear seesaw model can be mapped onto the parametrisation used in Ref. [57] in the case of degenerate heavy neutrinos by taking $\text{Re[z]}\rightarrow\theta/2$. Note that in that reference the Majorana phase is shifted by $\pi/2$ in order to provide all positive neutrino masses, whereas here the result is automatically obtained.
@@
III.2.2 Loop level: the Generalised Scotogenic Model
Here we discuss the Generalised Scotogenic Model (GScM) [53], a generalisation of the ScM Model [93, 72] based on a global $\text{U}(1)_{\text{DM}}$ symmetry instead of the $\mathbb{Z}_{2}$ symmetry of ScM. We follow closely the notation of Ref. [53]. The SM is augmented by two additional scalar doublets and at least one vector-like Dirac fermion, all charged under the $\text{U}(1)_{\text{DM}}$ symmetry. The particle content with the associated charges can be found in Table 2. The relevant terms in the Lagrangian are
@@ III.2.2 Loop level: the Generalised Scotogenic Model
$\mathcal{V}_{\rm GScM}$ $=-m_{H}^{2}H^{\dagger}H+\lambda_{H}(H^{\dagger}H)^{2}+m_{\Phi}^{2}\Phi^{\dagger}\Phi+\lambda_{\Phi}(\Phi^{\dagger}\Phi)^{2}$
$+m_{\Phi^{\prime}}^{2}\Phi^{\prime\dagger}\Phi^{\prime}+\lambda_{\Phi^{\prime}}(\Phi^{\prime\dagger}\Phi^{\prime})^{2}+\lambda_{H\Phi}(H^{\dagger}H)(\Phi^{\dagger}\Phi)$
$++\lambda_{H\Phi^{\prime}}(H^{\dagger}H)(\Phi^{\prime\dagger}\Phi^{\prime})+\lambda_{\Phi\Phi^{\prime}}(\Phi^{\dagger}\Phi)(\Phi^{\prime\dagger}\Phi^{\prime})$
$+\lambda_{H\Phi,2}(H^{\dagger}\Phi)(\Phi^{\dagger}H)+\lambda_{H\Phi^{\prime},2}(H^{\dagger}\tilde{\Phi}^{\prime})(\tilde{\Phi}^{\prime\dagger}H)$
$+\lambda_{\Phi\Phi^{\prime},2}(\Phi^{\dagger}\tilde{\Phi}^{\prime})(\tilde{\Phi}^{\prime\dagger}\Phi)+\lambda_{H\Phi\Phi^{\prime}}\left[(H^{\dagger}\tilde{\Phi}^{\prime})(H^{\dagger}\Phi)+\text{H.c.}\right].$ (43)
@@
$(m_{\nu})_{\alpha\beta}$ $=\dfrac{\sin 2\theta}{32\pi^{2}}\sum_{i}\left(Y_{\alpha i}Y^{\prime}_{\beta i}M_{i}\,+Y^{\prime}_{\alpha i}Y_{\beta i}M_{i}\,\right)$
$\times\mathcal{F}(m_{\eta_{0}},m_{\eta_{0}^{\prime}},M_{i})\,$, (49)
@@
$Y_{1}=Y,\,Y_{2}=Y^{\prime},\,M_{12}=\dfrac{\sin 2\theta}{32\pi^{2}}\,{\rm diag}(M_{i}\mathcal{F}(M_{i}))\,.$ (50)
For the $n_{1}=n_{2}=1$ case, Eqs. (III.2) and (III.2) may be used for IO and NO, respectively. As a consistency check, we note that the parametrisation derived from the GCI construction coincides with that employed in Refs. [22, 53] by replacing $\zeta\rightarrow e^{-\text{Im}z}$. For suitable lepton number charge assignments of the additional scalar and fermion fields, lepton number violation originates from the Yukawa coupling $Y$, which can thus be regarded as naturally small in the sense of ’t Hooft naturalness. This corresponds to the large imaginary part of the complex angle $z$.
@@
III.3 Diagonal plus off-diagonal mass matrix